OpenStudy (anonymous):

How do you solve this? (x to the 6th power y to the 4th power) -1\2?

OpenStudy (anonymous):

\[(x6 y ^4)-1\2\]

OpenStudy (anonymous):

Are you solving for x or y or are you solving for when they = 0

OpenStudy (anonymous):

I'm just trying to simply the equation and write the answer only using positive exponents

OpenStudy (campbell_st):

ok... use the powe of a power law \[(x^a)^b = x^{a \times b}\] a your problem can be split \[(x^6)^{\frac{-1}{2}} \times (y^4)^{\frac{-1}{2}}\] use the law above to simplify them

OpenStudy (anonymous):

so multiply 6 by -1\2 and 4 by -1\2?

OpenStudy (campbell_st):

then you will be able to write them as fractions.

OpenStudy (campbell_st):

yep... thats the start

OpenStudy (campbell_st):

then negative powers indicate fractions \[x^{-a} = \frac{1}{x^a}\]

OpenStudy (anonymous):

so then would it be (x ^3) (y ^2)

OpenStudy (campbell_st):

well the powers after multiplying are \[x^{-3}y^{-2}\]

OpenStudy (campbell_st):

now use the negative index law

OpenStudy (anonymous):

ya I realized they should've been negative

OpenStudy (anonymous):

so now do x ^-3 y ^-2

OpenStudy (anonymous):

and then x to the -3 power = 1 over x cubed and the same thing for y, but with 2 for that one?

OpenStudy (anonymous):

yes x^-3 you always want to flip to the bottom so 1/x^3 is the same thing but positive exponent.

OpenStudy (anonymous):

so now that I've done that, what's next?

OpenStudy (anonymous):

I guess according to him (1/x^3)*(1/y^2) you combine them so you should get something I think like 1/(X^3*y^2)

OpenStudy (anonymous):

That actually makes sense a little bit. Thanks for trying to help :)

OpenStudy (campbell_st):

thats correct...

OpenStudy (anonymous):

so I'm working another problem which is a ^3\4 and it needs to be turned into a radical. I have \[\sqrt a ^3 x \sqrt a ^4)\]

OpenStudy (anonymous):

am i on the right track?